# F-95000 Cergy-Pontoise, France

### v1

@@ 6.2 Analytic Queries with no Group-by Clause
33:  return  $proc\_change$, $proc\_min$, $proc\_max$, $proc\_ans^{*}$
Proposition 10
  • If $aggr$ is $min$, $max$ or $count$, then ${\sf C\_ans}(\mathcal{AQ})=[min\_ans,max\_ans]$ where $min\_ans$ and $max\_ans$ are returned by Algorithm 3.
  • If $aggr=sum$ and if ${\sf C\_ans}(\mathcal{AQ})=[glb,lub]$, then $min\_ans$ and $max\_ans$ as returned by Algorithm 3 satisfy that $min\_ans\leq glb$ and $max\_ans\geq lub$.
  • For every aggregate function and every selection condition $\Gamma$, $ans^{*}$ as returned by Algorithm 3 is equal to ${\sf C\_ans}^{*}(\mathcal{AQ})$.
Proof. See Appendix C. $\Box$
@@
0:  The m-table $m\_Chase(T)$ and $\mathcal{AQ}:~{}$select $aggr(M_{i})$ from $T$ where $\Gamma$
@@
$[min\_ans,max\_ans]$ // meant to be equal to ${\sf C\_ans}(\mathcal{AQ})$
@@
1:  Call procedure Compute_aggregate with input parameters $m\_Chase(T)$, $\mathcal{AQ}$ andwith output parameters $change\_min\_max$, $min\_ans$, $max\_ans$, $ans^{*}$
@@
8:           return  ($[0,0]$, 00)
In the following example we show that Algorithm 3 may fail to return exact values for $glb$ and $lub$ when the aggregate operator is $sum$, operating on positive values and negative values.
@@
Thus $m\_Chase(T)$ contains three m-tuples $\sigma_{1}$, $\sigma_{2}$, $\sigma_{3}$ defined over $U$ as follows:
[…]
and $T$ has four repairs denoted by $R_{i}$ for $i=1,\ldots,4$, whose sets $R^{\mathbf{K}}_{i}$ of tuples over $U$ are defined by:
$R^{\mathbf{K}}_{1}=\{(k_{1},k_{2},10,2,30),(k^{\prime}_{1},k^{\prime}_{2},-15,0,-10),(k_{1},k^{\prime}_{2},10,0,100)\}$
$R^{\mathbf{K}}_{2}=\{(k_{1},k_{2},10,2,30),(k^{\prime}_{1},k^{\prime}_{2},-15,3,-10),(k_{1},k^{\prime}_{2},10,3,100)\}$
$R^{\mathbf{K}}_{3}=\{(k_{1},k_{2},10,2,30),(k^{\prime}_{1},k^{\prime}_{2},20,0,-10),(k_{1},k^{\prime}_{2},10,0,100)\}$
$R^{\mathbf{K}}_{4}=\{(k_{1},k_{2},10,2,30),(k^{\prime}_{1},k^{\prime}_{2},20,3,-10),(k_{1},k^{\prime}_{2},10,3,100)\}$.
@@
    When running the call of the procedure Compute_aggregate in Algorithm 3, $min\_ans$ is first set to $30$ when processing $\sigma_{1}$, then to $\min(\{30,\min\{-10,30\}\})=-10$ when processing $\sigma_{2}$, and then to $\min(\{-10,\min\{100,-10\}\})=-10$ when processing $\sigma_{3}$. Similarly, $max\_ans$ is first set to $30$ when processing $\sigma_{1}$, then to $\max(\{30,\min\{-10,30\}\})=30$ when processing $\sigma_{2}$, and then to $\max(\{30,\min\{100,30\}\})=30$ when processing $\sigma_{3}$. Hence, Algorithm 3 returns $[-10,30]$ as expected.
@@
    When running the call of the procedure Compute_aggregate in Algorithm 3, $min\_ans$ is first set to $1$ when processing $\sigma_{1}$ and then unchanged when processing $\sigma_{2}$ and $\sigma_{3}$. Moreover, $max\_ans$ is increased by $1$ for each m-tuple $\sigma_{1}$, $\sigma_{2}$ and $\sigma_{3}$. Thus Algorithm 3 returns $[1,3]$ as expected.
@@
    When running the call of the procedure Compute_aggregate in Algorithm 3, $min\_ans$ is successively set to $30$, $(30-10)=20$ when processing $\sigma_{1}$ and then $\sigma_{2}$, and unchanged when processing $\sigma_{3}$. On the other hand, $max\_ans$ is successively set to $30$ when processing $\sigma_{1}$, unchanged when processing $\sigma_{2}$ and set to $(30+100)=130$ when processing $\sigma_{3}$. Thus the call of the procedure Compute_aggregate in Algorithm 3 returns $[20,130]$.
  • On the other hand, assuming that the second tuple in $F$ is $(k^{\prime}_{1}k^{\prime}_{2},10)$ (instead of $(k^{\prime}_{1}k^{\prime}_{2},-10)$), then ${\sf C\_ans}(\mathcal{AQ})=[30,140]$, because ${\sf Ans}(\mathcal{AQ}^{[R_{i}]})$ is $30$ if $i=1$ or $i=3$, $130$ if $i=2$ and $140$ if $i=4$. It can be seen that when running the call of the procedure Compute_aggregate in Algorithm 3, we obtain $min\_ans=30$ and $max\_ans=140$. $\Box$
@@
0:  The m-table $m\_Chase(T)$ and $\mathcal{AQ}:~{}$select $X,aggr(M_{i})$ from $T$ where $\Gamma$ group by $X$
@@
$Cons\_Ans$: a set of pairs $\langle x,[min\_ans,max\_ans]\rangle$ // meant to be equal to ${\sf C\_ans}(\mathcal{AQ})$
@@
2:  for all $\sigma$ in $m\_Chase(T)$ do
3:      if $\mathbf{K}\cup sch(Q)\subseteq sch(\sigma)$ then
4:           if $(Y\to B$ in $FD$ and $YB\subseteq X\Rightarrow|{\sf tuples}(\sigma(B))|=1)$ and $({\sf tuples}(\sigma(sch(\Gamma)))\cap Sat(\Gamma)\neq\emptyset$then
@@
6:  for all $X$-value $x$ occurring in $Temp$ do
@@
8:      Call procedure Compute_aggregate with input parameters $Temp(x)$, $\mathcal{AQ}$ andwith output parameters $change\_min\_max$, $min\_ans$, $max\_ans$, $ans^{*}$
@@ 6.3 Analytic Queries with a Group-by Clause
As in the case of analytic queries with no group-by clause, when the query involves a group-by clause, the computation of the consistent answers also involves a call of the procedure Compute_aggregate, as shown line 8 of Algorithm 4. This algorithm works as follows: a first scan of $m\_Chase(T)$, lines 2-5 retrieves in the set $Temp$ all relevant m-tuples $\sigma$, namely, such that $sch(\sigma)$ contains all attributes occurring in $\mathcal{AQ}$ and such that the $X$-values are consistent and may be associated with a tuple over $sch(\Gamma)$ satisfying the condition $\Gamma$. Then, a subsequent loop lines 6-11 operates a call of the procedure Compute_aggregate for each $X$-value $x$ occurring in the m-tuples of the set $Temp$. For each such call, Algorithm 2 scans the subset $Temp(x)$ of $Temp$ and computes the corresponding aggregate values appearing in the answers associated with $x$.
@@
  • If $aggr=min$, $aggr=max$ or $aggr=count$, then $Cons\_Ans$ as returned by Algorithm 4 is equal to ${\sf C\_ans}(\mathcal{AQ})$.
  • If $aggr=sum$, then $\langle x,[min\_ans,max\_ans]\rangle$ is in $Cons\_Ans$ as returned by Algorithm 4 if and only if ${\sf C\_ans}(\mathcal{AQ})$ contains $\langle x,[lub,glb]\rangle$ such that $min\_ans\leq glb$ and $max\_ans\geq lub$.
@@
  • For every aggregate function and every selection condition $\Gamma$, $Ans^{*}$ as returned by Algorithm 4 is equal to ${\sf C\_ans}^{*}(\mathcal{AQ})$.
Proof. First, the loop lines 2-5 scans $m\_Chase(T)$ and collects in the set $Temp$ the only m-tuples necessary to compute the consistent answer. Indeed, the collected m-tuples are all m-tuples $\sigma$ such that: they are defined over a super-set of $\mathbf{K}\cup sch(\sigma)$, their $X$-component contains consistent tuples, and their m-tuples over $sch(\Gamma)$ contain at least one tuple of $Sat(\Gamma)$. This is so because any m-tuple in $m\_Chase(T)$ not satisfying the above conditions can not contribute to the consistent answer. Then, in the next steps for every collected $X$-value, its associated aggregate value is evaluated as done in Algorithm 3. As a consequence of Proposition 10, for $x$ in $Temp$, the loop lines 6-11 in Algorithm 4 generates correct answers. The proof is therefore complete. $\Box$
We emphasize that Algorithm 3 and Algorithm 4 show that the consistent answers to a given analytic query $\mathcal{AQ}$, with or without a group-by clause, are computed in linear time in the size of $m\_Chase(T)$. Recalling that the computation of $m\_Chase(T)$ has been shown in LauS25 to be polynomial in time with respect to the size of $T$, it turns out that ${\sf C\_ans}(\mathcal{AQ})$ and ${\sf C\_ans}^{*}(\mathcal{AQ})$ can be computed in polynomial time with respect to the size of $T$.
@@ 6.4 Analytic Queries with a group-by-having Clause
  • If we assume that $m_{1}=5$, $m^{\prime}_{1}=5$ and $m^{\prime\prime}_{1}=15$, then ${\sf C\_ans}(\mathcal{AQ}^{\prime}_{3})=\{\langle a_{2},[10,20],[5,15]\rangle\}$ and so, ${\sf C\_ans}(\mathcal{AQ}_{3})$ is empty. Notice incidentally that it is easily seen from Figure 3 that there exist repairs $R$ in ${\sf Rep}(T)$ for which ${\sf Ans}(\mathcal{AQ}_{3}^{{}^{\prime}[R]})=\{\langle a_{2},20,15\rangle\}$, implying that ${\sf Ans}(\mathcal{AQ}_{3}^{[R]})=\emptyset$.
  • If we assume that $m_{1}=5$, $m^{\prime}_{1}=5$ and $m^{\prime\prime}_{1}=8$, then ${\sf C\_ans}(\mathcal{AQ}^{\prime}_{3})=\{\langle a_{2},[10,13],[5,8]\rangle\}$ and so, $\langle a_{2},[10,13]\rangle$ is in ${\sf C\_ans}(\mathcal{AQ}_{3})$. In this case, it is easily seen from Figure 3 that for every repair $R$ in ${\sf Rep}(T)$, ${\sf Ans}(\mathcal{AQ}_{3}^{{}^{\prime}[R]})$ contains a pair $\langle a_{2},\mu,\nu\rangle$ where $10\leq\mu\leq 13$ and $\nu<10$, implying that $\langle a_{2},[10,13]\rangle$ is in ${\sf C\_ans}(\mathcal{AQ}_{3})$.
@@
[…]
@@
such that the above restrictions are met. If $\Upsilon=\Upsilon_{1}\wedge\ldots\wedge\Upsilon_{h}$, let $\mathcal{AQ}^{\prime}$ be the following analytic query with no having clause:
@@
Then, $\langle x,[glb,lub]\rangle$ belongs to ${\sf C\_ans}(\mathcal{AQ})$ if and only if there exists $\langle x,[glb,lub],$ $[glb_{1},lub_{1}],\ldots,[glb_{h},lub_{h}]\rangle$ in ${\sf C\_ans}(\mathcal{AQ}^{\prime})$ such that for every $p=1,\ldots,h$, $[glb_{p},lub_{p}]\subseteq Sat(\Upsilon_{p})$.
Proof. Let $\langle x,[glb,lub],$ $[glb_{1},lub_{1}],\ldots,[glb_{h},lub_{h}]\rangle$ in ${\sf C\_ans}(\mathcal{AQ}^{\prime})$ such that for every $p=1,\ldots,h$, $[glb_{p},lub_{p}]\subseteq Sat(\Upsilon_{p})$. Given $R$ in ${\sf Rep}(T)$, ${\sf Ans}(\mathcal{AQ}^{{}^{\prime}[R]})$ contains a tuple $\langle x,\mu,\mu_{1},\ldots,\mu_{p}\rangle$ such that $\mu\in[glb,lub]$ and $\mu_{p}\in[glb_{p},lub_{p}]$ for $p=1,\ldots,h$ hold. Hence, for $p=1,\ldots,h$, $\mu_{p}$ is in $Sat(\Upsilon_{p})$, implying that $\langle x,[glb,lub]\rangle$ is in ${\sf C\_ans}(\mathcal{AQ})$.
Conversely, let $\langle x,[glb,lub]\rangle$ be in ${\sf C\_ans}(\mathcal{AQ})$ and let $p$ such that $[glb_{p},lub_{p}]\subseteq Sat(\Upsilon_{p})$ does not hold. Since $Sat(\Upsilon_{p})$ is an interval, this implies that $glb_{p}\not\in Sat(\Upsilon_{p})$ or that $lub_{p}\not\in Sat(\Upsilon_{p})$. Since there exists a repair $R_{min}$, respectively a repair $R_{max}$, such that $aggr_{p}(M_{i_{p}})=glb_{p}$, respectively $aggr_{p}(M_{i_{p}})=lub_{p}$, it turns out that there exists at least one repair $R$ such that $ans(\mathcal{AQ}^{{}^{\prime}[R]})$ contains a tuple $\langle x,\mu,\mu_{1},\ldots,\mu_{h}\rangle$ such that $\mu_{p}\not\in Sat(\Upsilon_{p})$. Thus $\langle x,\mu\rangle$ is not in $ans(\mathcal{AQ}^{[R]})$, which implies that $\langle x,[glb,lub]\rangle$ is not in ${\sf C\_ans}(\mathcal{AQ})$; a contradiction which completes the proof. $\Box$
It should be clear from Proposition 11 and Proposition 12 that, under the restrictions earlier stated, the consistent answer of an analytic query involving a group-by-having clause can be computed in polynomial time with respect to the size of $T$.

### v2

@@ 6.1 Basic Result
33:  return proc_changeproc\_change, proc_minproc\_min, proc_maxproc\_max, $proc\_ans^{*}$
In the following two sections 6.2 and 6.3, we successively introduce our algorithms for computing consistent answers to analytic queries, depending on whether they involve or not a group-by clause. In each case, the main algorithm relies on a procedure, called Compute_aggregate and shown in Algorithm 3, whose role is to scan (either entirely or partially) the m-table m_Chase(T)m\_Chase(T) and return values that will appear in the answers.
6.2 Analytic Queries with no Group-by Clause
If the query involves no group-by clause, as shown in Algorithm 4, the procedure Compute_aggregate of Algorithm 3 is called to scan the whole m-table m_Chase(T)m\_Chase(T). When running this call, line 1 in Algorithm 4, the main loop lines 6-32 of Algorithm 3 scans m_Chase(T)m\_Chase(T) and computes values appearing in the answers. The following proposition shows that Algorithm 4 is correct, except when $aggr=sum$. In other words, if $aggr\neq sum$ Algorithm 4 returns ${\sf C\_ans}(\mathcal{AQ})$. Moreover, it is also shown that ${\sf C\_ans}^{*}(\mathcal{AQ})$ is correctly computed by Algorithm 4 in any case.
@@ 6.2 Analytic Queries with no Group-by Clause
0:  The m-table m_Chase(T)m\_Chase(T) and $\mathcal{AQ}:$~select $aggr(M_{i})$ from $T$ where $\Gamma$
@@
[min_ans,max_ans][min\_ans,max\_ans] // meant to be equal to ${\sf C\_ans}(\mathcal{AQ})$
@@
1:  Call procedure Compute_aggregate with input parameters m_Chase(T)m\_Chase(T), $\mathcal{AQ}$ andwith output parameters change_min_maxchange\_min\_max, min_ansmin\_ans, max_ansmax\_ans, $ans^{*}$
@@
8:     return ($[0,0]$, 0)
Proposition 11
  • If aggraggr is minmin, maxmax or countcount, then ${\sf C\_ans}(\mathcal{AQ})=[min\_ans,max\_ans]$ where min_ansmin\_ans and max_ansmax\_ans are returned by Algorithm 4.
  • If $aggr=sum$ and if ${\sf C\_ans}(\mathcal{AQ})=[glb,lub]$, then min_ansmin\_ans and max_ansmax\_ans as returned by Algorithm 4 satisfy that $min\_ans\leq glb$ and $max\_ans\geq lub$.
  • For every aggregate operator and every selection condition $\Gamma$, $ans^{*}$ as returned by Algorithm 4 is equal to ${\sf C\_ans}^{*}(\mathcal{AQ})$.
Proof. See Appendix G. $\Box$
In the following example we show that Algorithm 4 may fail to return exact values for glbglb and lublub when the aggregate operator is sumsum, operating on positive values and negative values.
@@
Thus m_Chase(T)m\_Chase(T) contains three m-tuples $\sigma_{1}$, $\sigma_{2}$, $\sigma_{3}$ defined over $U$ as follows:
$\sigma_{1}=(k_{1})(k_{2})(10)(2)(30)$, $\sigma_{2}=(k^{\prime}_{1})(k^{\prime}_{2})(-15\,20)(0\,3)(-10)$, $\sigma_{3}=(k_{1})(k^{\prime}_{2})(10)(0\,3)(100)$
and $T$ has four repairs denoted by $R_{i}$ for $i=1,\ldots,4$, whose sets $R^{\mathbb{K}}_{i}$ of tuples over $U$ are defined by:
$R^{\mathbb{K}}_{1}=\{(k_{1},k_{2},10,2,30),(k^{\prime}_{1},k^{\prime}_{2},-15,0,-10),(k_{1},k^{\prime}_{2},10,0,100)\}$
$R^{\mathbb{K}}_{2}=\{(k_{1},k_{2},10,2,30),(k^{\prime}_{1},k^{\prime}_{2},-15,3,-10),(k_{1},k^{\prime}_{2},10,3,100)\}$
$R^{\mathbb{K}}_{3}=\{(k_{1},k_{2},10,2,30),(k^{\prime}_{1},k^{\prime}_{2},20,0,-10),(k_{1},k^{\prime}_{2},10,0,100)\}$
$R^{\mathbb{K}}_{4}=\{(k_{1},k_{2},10,2,30),(k^{\prime}_{1},k^{\prime}_{2},20,3,-10),(k_{1},k^{\prime}_{2},10,3,100)\}$.
@@
    When running the call of the procedure Compute_aggregate in Algorithm 4, min_ansmin\_ans is first set to $30$ when processing $\sigma_{1}$, then to $\min(\{30,\min\{-10,30\}\})=-10$ when processing $\sigma_{2}$, and then to $\min(\{-10,\min\{100,-10\}\})=-10$ when processing $\sigma_{3}$. Similarly, max_ansmax\_ans is first set to $30$ when processing $\sigma_{1}$, then to $\max(\{30,\min\{-10,30\}\})=30$ when processing $\sigma_{2}$, and then to $\max(\{30,\min\{100,30\}\})=30$ when processing $\sigma_{3}$. Hence, Algorithm 4 returns [10,30][-10,30] as expected.
@@
    When running the call of the procedure Compute_aggregate in Algorithm 4, min_ansmin\_ans is first set to $1$ when processing $\sigma_{1}$ and then unchanged when processing $\sigma_{2}$ and $\sigma_{3}$. Moreover, max_ansmax\_ans is increased by $1$ for each m-tuple $\sigma_{1}$, $\sigma_{2}$ and $\sigma_{3}$. Thus Algorithm 4 returns $[1,3]$ as expected.
@@
    When running the call of the procedure Compute_aggregate in Algorithm 4, min_ansmin\_ans is successively set to $30$, (3010)=20(30-10)=20 when processing $\sigma_{1}$ and then $\sigma_{2}$, and unchanged when processing $\sigma_{3}$. On the other hand, max_ansmax\_ans is successively set to $30$ when processing $\sigma_{1}$, unchanged when processing $\sigma_{2}$ and set to $(30+100)=130$ when processing $\sigma_{3}$. Thus the call of the procedure Compute_aggregate in Algorithm 4 returns $[20,130]$.
  • On the other hand, assuming that the second tuple in $F$ is $(k^{\prime}_{1}k^{\prime}_{2},10)$ (instead of $(k^{\prime}_{1}k^{\prime}_{2},-10)$), then ${\sf C\_ans}(\mathcal{AQ})=[30,140]$, because ${\sf Ans}(\mathcal{AQ}^{[R_{i}]})$ is $30$ if $i=1$ or $i=3$, $130$ if $i=2$ and $140$ if $i=4$. It can be seen that when running the call of the procedure Compute_aggregate in Algorithm 4, we obtain min_ans=30min\_ans=30 and max_ans=140max\_ans=140. $\Box$
@@
0:  The m-table m_Chase(T)m\_Chase(T) and $\mathcal{AQ}:$~select $X,aggr(M_{i})$ from $T$ where $\Gamma$ group by $X$
@@
Cons_AnsCons\_Ans: a set of pairs $\langle x,[min\_ans,max\_ans]\rangle$ // meant to be equal to ${\sf C\_ans}(\mathcal{AQ})$
@@
2:  for all $\sigma$ in m_Chase(T)m\_Chase(T) do
3:    if $\mathbb{K}\cup sch(Q)\subseteq sch(\sigma)$ then
4:     if $(Y\to B$ in $FD$ and $YB\subseteq X\Rightarrow|{\sf tuples}(\sigma(B))|=1)$ and $({\sf tuples}(\sigma(sch(\Gamma)))\cap Sat(\Gamma)\neq\emptyset$) then
@@
6:  for all $X$-value $x$ occurring in TempTemp do
@@
8:    Call procedure Compute_aggregate with input parameters $Temp(x)$, $\mathcal{AQ}$ andwith output parameters change_min_maxchange\_min\_max, min_ansmin\_ans, max_ansmax\_ans, $ans^{*}$
@@ 6.3 Analytic Queries with a Group-by Clause
As in the case of analytic queries with no group-by clause, when the query involves a group-by clause, the computation of the consistent answers also involves a call of the procedure Compute_aggregate, as shown line 8 of Algorithm 5. This algorithm works as follows: a first scan of m_Chase(T)m\_Chase(T), lines 2-5 retrieves in the set $Temp$ all relevant m-tuples $\sigma$, namely, such that $sch(\sigma)$ contains all attributes occurring in $\mathcal{AQ}$ and such that the $X$-values are consistent and may be associated with a tuple over $sch(\Gamma)$ satisfying the condition $\Gamma$. Then, a subsequent loop lines 6-11 operates a call of the procedure Compute_aggregate for each $X$-value $x$ occurring in the m-tuples of the set $Temp$. For each such call, Algorithm 3 scans the subset $Temp(x)$ of $Temp$ and computes the corresponding aggregate values appearing in the answers associated with $x$.
@@
  • If $aggr=min$, $aggr=max$ or $aggr=count$, then Cons_AnsCons\_Ans as returned by Algorithm 5 is equal to ${\sf C\_ans}(\mathcal{AQ})$.
  • If $aggr=sum$, then $\langle x,[min\_ans,max\_ans]\rangle$ is in Cons_AnsCons\_Ans as returned by Algorithm 5 if and only if ${\sf C\_ans}(\mathcal{AQ})$ contains $\langle x,[lub,glb]\rangle$ such that $min\_ans\leq glb$ and $max\_ans\geq lub$.
@@
  • For every aggregate operator and every selection condition $\Gamma$, $Ans^{*}$ as returned by Algorithm 5 is equal to ${\sf C\_ans}^{*}(\mathcal{AQ})$.
Proof. First, the loop lines 2-5 scans m_Chase(T)m\_Chase(T) and collects in the set $Temp$ the only m-tuples necessary to compute the consistent answer. Indeed, the collected m-tuples are all m-tuples $\sigma$ such that: they are defined over a super-set of $\mathbb{K}\cup sch(\sigma)$, their $X$-component contains consistent tuples, and their m-tuples over $sch(\Gamma)$ contain at least one tuple of $Sat(\Gamma)$. This is so because any m-tuple in m_Chase(T)m\_Chase(T) not satisfying the above conditions can not contribute to the consistent answer. Then, in the next steps for every collected $X$-value, its associated aggregate value is evaluated as done in Algorithm 4. As a consequence of Proposition 11, for $x$ in $Temp$, the loop lines 6-11 in Algorithm 5 generates correct answers. The proof is therefore complete. $\Box$
We emphasize that Algorithm 4 and Algorithm 5 show that the consistent answers to a given analytic query $\mathcal{AQ}$, with or without a group-by clause, are computed in linear time in the size of m_Chase(T)m\_Chase(T). Recalling that the computation of m_Chase(T)m\_Chase(T) has been shown earlier to be quasi-linear with respect to the size of $T$ (or in $\mathcal{O}(W\cdot\log(W))$ where $W$ is the size of the data warehouse), it turns out that ${\sf C\_ans}(\mathcal{AQ})$ and ${\sf C\_ans}^{*}(\mathcal{AQ})$ can be computed in quasi-linear time with respect to the size of the data warehouse.
@@ 6.4 Analytic Queries with a group-by-having Clause
  • If we assume that $m_{1}=5$, $m^{\prime}_{1}=5$ and $m^{\prime\prime}_{1}=15$, then ${\sf C\_ans}(\mathcal{AQ}^{\prime}_{3})=\{\langle a_{2},[10,20],[5,15]\rangle\}$ and so, ${\sf C\_ans}(\mathcal{AQ}_{3})$ is empty. Notice incidentally that it is easily seen from Figure 4 that there exist repairs $R$ in ${\sf Rep}(T)$ for which ${\sf Ans}(\mathcal{AQ}_{3}^{\prime[R]})=\{\langle a_{2},20,15\rangle\}$, implying that ${\sf Ans}(\mathcal{AQ}_{3}^{[R]})=\emptyset$.
  • If we assume that $m_{1}=5$, $m^{\prime}_{1}=5$ and $m^{\prime\prime}_{1}=8$, then ${\sf C\_ans}(\mathcal{AQ}^{\prime}_{3})=\{\langle a_{2},[10,13],[5,8]\rangle\}$ and so, $\langle a_{2},[10,13]\rangle$ is in ${\sf C\_ans}(\mathcal{AQ}_{3})$. In this case, it is easily seen from Figure 4 that for every repair $R$ in ${\sf Rep}(T)$, ${\sf Ans}(\mathcal{AQ}_{3}^{\prime[R]})$ contains a pair $\langle a_{2},\mu,\nu\rangle$ where $10\leq\mu\leq 13$ and $\nu<10$, implying that $\langle a_{2},[10,13]\rangle$ is in ${\sf C\_ans}(\mathcal{AQ}_{3})$.
@@
    $\Upsilon$ is a conjunctive boolean expression built up from atoms of the form $aggr(M_{i})\,\theta\,\alpha$ where $M_{i}$ is a measure attribute in $\mathbb{M}$, $\theta$ is a comparison predicate in ${<,,>,}\{$$\leq$$,\geq\}$ and $\alpha$ is a number.
@@
such that the above two restrictions regarding $\Upsilon$ are met. If $\Upsilon=\Upsilon_{1}\wedge\ldots\wedge\Upsilon_{h}$, let $\mathcal{AQ}^{\prime}$ be the following analytic query with no having clause:
@@
The tuple $\langle x,[glb,lub]\rangle$ belongs to ${\sf C\_ans}(\mathcal{AQ})$ if and only if there exists a tuple of the form $\langle x,[glb,lub],$ $glb_{1},lub_{1}],\ldots,[glb_{h},lub_{h}]\rangle$ in ${\sf C\_ans}(\mathcal{AQ}^{\prime})$ such that $glb_{p},lub_{p}]\subseteq Sat(\Upsilon_{p})$ holds for every $p=1,\ldots,h$.
Proof. Let $\langle x,[glb,lub],$ $glb_{1},lub_{1}],\ldots,[glb_{h},lub_{h}]\rangle$ in ${\sf C\_ans}(\mathcal{AQ}^{\prime})$ such that for every $p=1,\ldots,h$, $glb_{p},lub_{p}]\subseteq Sat(\Upsilon_{p})$. Given $R$ in ${\sf Rep}(T)$, ${\sf Ans}(\mathcal{AQ}^{\prime[R]})$ contains a tuple $\langle x,\mu,\mu_{1},\ldots,\mu_{p}\rangle$ such that $\mu\in[glb,lub]$ and $\mu_{p}\in[glb_{p},lub_{p}]$ for $p=1,\ldots,h$ hold. Hence, for $p=1,\ldots,h$, $\mu_{p}$ is in $Sat(\Upsilon_{p})$, implying that $\langle x,[glb,lub]\rangle$ is in ${\sf C\_ans}(\mathcal{AQ})$.
Conversely, let $\langle x,[glb,lub]\rangle$ be in ${\sf C\_ans}(\mathcal{AQ})$ and let $p$ such that $glb_{p},lub_{p}]\subseteq Sat(\Upsilon_{p})$ does not hold. Since $Sat(\Upsilon_{p})$ is an interval, this implies that $glb_{p}\not\in Sat(\Upsilon_{p})$ or that $lub_{p}\not\in Sat(\Upsilon_{p})$. Since there exists a repair $R_{min}$, respectively a repair $R_{max}$, such that $aggr_{p}(M_{i_{p}})=glb_{p}$, respectively $aggr_{p}(M_{i_{p}})=lub_{p}$, it turns out that there exists at least one repair $R$ such that $ans(\mathcal{AQ}^{\prime[R]})$ contains a tuple $\langle x,\mu,\mu_{1},\ldots,\mu_{h}\rangle$ such that $\mu_{p}\not\in Sat(\Upsilon_{p})$. Thus $\langle x,\mu\rangle$ is not in $ans(\mathcal{AQ}^{[R]})$, which implies that $\langle x,[glb,lub]\rangle$ is not in ${\sf C\_ans}(\mathcal{AQ})$; a contradiction which completes the proof. $\Box$
It should be clear from Proposition 12 and Proposition 13 that, under the restrictions regarding $\Upsilon$ earlier stated, the consistent answer of an analytic query involving a group-by-having clause can be computed in quasi-linear time with respect to the size of the data warehouse.